# Generalized Bayesian deep reinforcement learning

### v1

@@ 3.3 Sequential Monte Carlo with gradient based kernel
     Evaluate $\nabla\tilde{U}(\bm{\theta}_{(t-1)})$
     $\mathbf{p}_{(t)}=\mathbf{p}_{(t-1)}-(\bm{\eta}_{(t-1)}\mathbf{p}_{(t-1)}+\nabla\tilde{U}(\bm{\theta}_{(t-1)})G(\bm{\theta}_{(t-1)}))\epsilon+\mathcal{N}(0,2a\epsilon G(\bm{\theta}_{(t-1)}))$
     $\bm{\theta}_{(t)}=\bm{\theta}_{(t-1)}+\mathbf{p}_{(t)}G(\bm{\theta}_{(t-1)})\epsilon$
     $\bm{\eta}_{(t)}=\bm{\eta}_{(t-1)}+(\frac{1}{d}\mathbf{p}_{(t)}^{T}\mathbf{p}_{(t)}-1)\epsilon$
@@
Note that, for our implementation, we have calculated the preconditioning matrix $G(\theta)$ using the gradient estimate as described in Chen et al. (2016).
@@ 4 ETS with generalized posterior
The above theorem suggests, irrespective of the number of episodes observed, the sequence of policies obtained from a policy iteration method integrated with ETS converges progressively toward optimal behavior as the iterations increase. The second term $\zeta_{l}(k,n)$ involves the samples drawn from the posterior of the model parameters obtained after observing up to $k$ episodes of interaction with the environment. As the episode count $k$ and the number of posterior samples $n$ increase, this term shrinks to 00 for any $l=1,2,\ldots$ according to Theorem 2 as the posterior distribution of the model parameters concentrates around the Expected SR minimizer. The detailed proof of the above theorem can be found in Appendix A.2.
@@ 5.1 Finite action MDP
To demonstrate the Scoring rule Thompson Sampling using SMC, we use the ‘inverted pendulum’ task from Dimitrakakis and Tziortziotis (2013). The agent targets to keep the pendulum upright as long as possible by switching actions. State of the environment is defined as (angle, angular velocity) of the pendulum. There are three possible actions from each state. The actions are force (in Newtons) in a certain direction, the action space is $(+50,0,-50)$. The problem has 6 parameters: the pendulum mass, the cart mass, the pendulum length, the gravity, the amount of uniform noise, and the simulation time interval. In this environment, agent gets $+1$ reward for every balancing step. The actual environment parameters are $\theta_{0}=(2.0,8.0,0.5,9.8,10,0.01)$ and the simulator model ($P_{\theta}$) for the dynamics of the environment is known to us.
In our comparison, we evaluated the online performance of both the model-free and Bayesian model-based approaches in the pendulum domain. In the model-free approach, we updated our policy after each episode of interaction with the environment. The policy updates ceased once we reached the maximum reward, which is known to be $1000$ in our setup.
@@
For policy learning, we adopted the least squares policy iteration (LSPI) method proposed by Lagoudakis and Parr (2003). LSPI is a model-free approximate policy iteration method that leverages the least squares temporal-difference learning algorithm to approximate the $Q$ function. Specifically, we approximated the $Q$ function using a linear combination of a $4\times 4$ grid of Gaussian radial basis functions in LSPI, with a learning rate of $\gamma=0.99$.
Additionally, for the Bayesian approach, we assumed a uniform prior over the interval $[0.5\theta_{0},5.0\theta_{0}]$ for the model parameters. We obtained samples from the posterior distribution after each episode using SMC with the adSGLD kernel (Jones and Leimkuhler, 2011) (refer to Table 1 for details of the hyperparameters). To define the intermediate target distributions between consecutive episodes, we chose the sequence of temperatures $\{\alpha_{k,l}\}_{l=1}^{L}$ at the $k$-th episode, such that the effective sample size (ESS) declines uniformly through out the SMC iterations. We used bisection method to find the temperature, $\alpha_{k,l}$ such that $\text{ESS}_{l}=\kappa\times\text{ESS}_{l-1}$, the parameter $\kappa$ for each experiment is mentioned in Appendix A.3. Once the ESS drops below half the original sample size, the particles are resampled. Here we used the zeroth order gradient (discussed in A.4) to compute an estimate of the gradient of the prequential loss. We used LSPI integrated with ETS to learn an optimal policy, as described in Section 2.1 where we updated the policy for a maximum of $J=30$ iterations in-between each episode.
Figure 4 illustrates that both the model-free and Bayesian model-based methods exhibit improved policy learning with increased training data. However, the Bayesian approach achieves optimal rewards much sooner compared to model-free training. This expedited convergence in Bayesian methods can be attributed to the fact that once the posterior distribution of the model parameters converges, simulated interactions closely resemble true interactions. With this concentrated posterior distribution, longer trajectories of simulated data can be generated, facilitating more accurate estimation of the value function. Consequently, when using these longer chains of simulated data to estimate the value function and perform LSPI to find an optimal policy, convergence occurs more rapidly due to the utilization of a more stable estimate of the value function.
@@
To demonstrate the application of ETS for the case where a simulator model is not known, we revisit the ‘inverted pendulum’ experiment with the previously mentioned parameters. Conditional GANs (Charlesworth and Montana, 2020; Zhao et al., 2021) have been widely adopted in the literature to model the dynamics. We have used a generative neural network, skipping the adversarial training, to model the difference between the next state and the current state conditioned on the current state and action as suggested by Nagabandi et al. (2018) and Deisenroth et al. (2013). Further, for the states $s$ that represent the angle, we have considered feeding in $(\sin(s),\cos(s))$ as inputs to the model. The model can be written as,
@@
For parameter inference we have used Sequential Monte Carlo (SMC) (implementation details can be found in Section A.3 in the Appendix). In high-dimensional parameter spaces, selecting a well-chosen initial sample from an informative prior is crucial for effective Bayesian inference. To do so, we ran the Adam optimizer (Kingma, 2014) with a learning rate of $0.001$ for $1000$ steps, saving the final $100$ points. The covariance of these optimized points was computed and used to add Gaussian noise, generating a total of $300$ particles to initialize the SMC. To define the intermediate target distributions between consecutive episodes, we chose the sequence of temperatures $\{\alpha_{k,l}\}_{l=1}^{L}$ at the $k$-th episode, such that the conditional effective sample size (CESS) (Zhou et al., 2016) stays constant through out the SMC iterations. We used bisection method to find the temperature, $\alpha_{k,l}$ such that $\text{ESS}_{l}=\kappa\times N$, where $N$ is the sample size and the parameter $\kappa$ for each experiment is mentioned in Appendix A.3. Once the ESS drops below $N/2$, the particles are resampled.
@@
Figure 5 demonstrates that LSPI combined with ETS learns the optimal policy significantly faster than its model-free counterpart. Although the true dynamics of the environment were unknown, the generative neural network effectively learned the dynamics, accelerating the policy learning process. Note that the Theorem 2 on the consistency of the generalized posterior distribution and the Theorem 3 on the bound on the approximation error of the $Q$ function under ETS assume a unimodal posterior for model parameters. Whilst this assumption may not hold when using a generative neural network as a simulator model, we still demonstrate that ETS is more sample-efficient for simpler models.
@@ 5.2 Continuous action MDP
When the action space is continuous, calculating the $Q$-function for all possible state-action pairs is impossible. A deterministic policy that maximizes the $Q$-function over the entire action space cannot be easily found. In such cases, a parameterized probabilistic policy is used, where $\mu_{\alpha}(a|s)$ denotes the probability of taking action $a$ when the environment is in state $s$, and $\alpha\in\mathbb{R}^{d^{\prime}}$ are the policy parameters. Usually the policy distribution is considered as a Gaussian distribution and $\alpha$ would contain the mean and covariance of the distribution. The agent here tries to update the policy parameters according based on the rewards collected during realtime interaction with the environment.
@@
To demonstrate the ETS strategy on a problem with continuous action space for the case of a misspecified model, we choose the ‘Hopper’ experiment from the OpenAI Gymnasium package (Kwiatkowski et al., 2024). This problem involves moving a two-dimensional, single-legged structure forward by applying force to three joints connecting its four body parts. Actions here consist of torques (ranging from -1 Nm to 1 Nm) applied to each of the three joints, while the environment’s state is defined as a $12$-dimensional real-valued vector of joint angles, angular velocities, positions, and velocities of the body parts. The reward function incentivizes forward movement of the hopper while penalizing the application of excessive torque, which could destabilize the system.
We have used a similar neural network architecture as described in Sction 5.1, for our simulator model, using three fully connected layers with $20$ nodes each. For sampling from this posterior of the model parameters, we used SMC, as before, but to maintain computational efficiency we used only $45$ particles (this figure was chosen based on the runs with 100 and 500 particles, which were not providing any significant improvements while already being expensive) for this problem. As before, we ran the Adam optimizer (Kingma, 2014) with a learning rate of $0.001$ for $1000$ steps, saving the final $15$ points. The covariance of these optimized points was computed and used to add Gaussian noise, generating a total of $45$ particles to initialize the SMC. Similar to misspecified ‘inverted pendulum’ problem, we defined intermediate target distributions for SMC based on the CESS.
For policy learning, we applied the ‘REINFORCE’ (Williams, 1992) a foundational policy gradient method suitable for continuous action spaces, integrating ETS via equation (15). The policy parameters were then updated based on equation (16). Full implementation details of the experiment can be found in Section A.3 and A.5 in the Appendix. Given that with an increasing amount of data, the posterior distribution can become overly concentrated, we limited posterior updates and sampling to the first $15$ episodes. At episode $15$, we halted model training and proceeded with the standard REINFORCE updates to refine the policy during future interactions with the environment.
Figure 6: For visual clarity we have considered the moving average (lag 10) of episodic returns collected using model-free REINFORCE and ETS-integrated REINFORCE. The average of the returns over five different seeds are plotted for both cases.
@@ 6 Conclusion
In this work, we have introduced a robust Bayesian framework for model-based reinforcement learning. A fully Bayesian treatment of model parameters traditionally requires a tractable likelihood function. This is often unavailable when the environment’s dynamics is learned through generative neural networks. To address this issue, we formulated a generalized posterior for the model parameters using prequential scoring rules based on a Markovian assumption on observed trajectories, enabling generalized Bayesian inference in the absence of a known likelihood. Additionally, we established a Bernstein-von Mises (BvM) type consistency result for the scoring rule-based generalized posterior in the discrete action setting, leaving the proof for continuous action space as future work.
@@
For efficient sampling from this generalized posterior, we used SMC samplers with an adSGLD kernel to leverage gradient information for better exploration in high-dimensional parameter spaces. We further improve the kernel with gradient-based preconditioning matrices similar to those in the Adam optimizer to provide better guidance for sampling in high-dimensions. For policy learning, we extended classical Thompson sampling by incorporating full posterior samples for enhanced policy search, introducing the expected Thompson sampling (ETS) approach. Using the BvM result, we derived an error bound for approximating the action-value function when a policy iteration method is integrated with ETS in well-specified model settings.
@@
To empirically evaluate the proposed method, we first compared classical model-free least squares policy iteration with the ETS-integrated Bayesian version on a simple inverted pendulum balancing task. In both well-specified and misspecified model settings, the ETS-integrated approach learned the optimal policy more quickly than the model-free baseline. We then applied ETS to a more complex problem of teaching a single-legged hopper to move forward without falling. Even with continuous action space and a misspecified model, the ETS-based approach discovered a high-reward policy significantly faster than the model-free approach.
@@
In conclusion, we present a robust framework for policy learning that enables rapid identification of high-reward policies in complex tasks. This approach is especially relevant for applications in robotics and autonomous systems where agents must make quick decisions without the luxury of prolonged learning periods. However, we observed that using generative deep neural networks as a simulator model leads to a multimodal posterior for model parameters. This makes sampling difficult, especially in high dimensions. We are keen to explore if a shrinkage prior on the parameter space can address this issue. Besides, future directions of this work include investigating copula-based nonparametric methods for model inference, potentially mitigating the effects of model misspecification. Another promising avenue would be integrating Bayesian model inference and policy training under a unified objective, optimizing both processes concurrently.
Appendix A
A.1 Proof of BvM, Theorem 2
Here we extend the proof of Bernstein–von Mises (BvM) theorem for generalized posterior (Miller, 2021) for our setup.
Proof  Let us define a function $g_{T}(x):\Theta\rightarrow\mathbb{R}$ as following,
$g_{T}(x)$ $=\exp\left(-\left[\operatorname{\mathcal{PS}}_{T}\left(\theta_{T}+x/\sqrt{T}\right)-\operatorname{\mathcal{PS}}_{T}\left(\theta_{T}\right)\right]\right)\pi\left(\theta_{T}+x/\sqrt{T}\right)$
$=q_{T}(x)z_{T}e^{\operatorname{\mathcal{PS}}_{T}\left(\theta_{T}\right)}T^{D/2}\quad\text{where, $q_{T}(x)=\pi_{\operatorname{\mathcal{PS}}}\left(\theta_{T}+x/\sqrt{T}|H_{T}\right)T^{-D/2}$ };$ (17)
recalling that $z_{T}<\infty$ by assumption, and define
$g_{0}(x)=\exp\left(-\frac{1}{2}x^{\prime}\operatorname{\Gamma}_{*}x\right)\pi\left(\theta^{*}\right)\text{with $\operatorname{\Gamma}_{*}=\sum_{a}w_{a}\operatorname{\Gamma}_{*}^{a}$}$
If $\lambda^{a}$ is the smallest eigenvalue of $H_{*}^{a}$ for each action $a\in\mathcal{A}$, then let us consider $\alpha^{a}\in(0,\lambda^{a})$ and a $\epsilon^{*}>0$ small enough such that $\epsilon^{*}<\alpha^{a}/\left(2c_{0}\right)\ \forall a\in\mathcal{A},\epsilon^{*}<\epsilon_{0}$, and $\pi(\theta)\leq 2\pi\left(\theta^{*}\right)$ for all $\theta\in B_{2\epsilon^{*}}\left(\theta^{*}\right)$ (as $\pi$ is continuous at $\theta^{*}$ ). As $\liminf_{T}\inf_{\theta\in B_{\epsilon^{*}}\left(\theta_{T}\right)^{c}}\frac{1}{|T_{a}|}(\operatorname{\mathcal{PS}}_{T}^{a}(\theta)-\operatorname{\mathcal{PS}}_{T}^{a}\left(\theta_{T}\right))>0$ for any $\epsilon>0$ from assumption A5, we define $\delta_{a}=\liminf_{T}\inf_{\theta\in B_{\epsilon^{*}}\left(\theta_{T}\right)^{c}}\frac{1}{|T_{a}|}(\operatorname{\mathcal{PS}}_{T}^{a}(\theta)-\operatorname{\mathcal{PS}}_{T}^{a}\left(\theta_{T}\right))>0,\forall a\in\mathcal{A}$ and $\delta=\sum_{a\in\mathcal{A}}\frac{|T_{a}|}{T}\delta_{a}.$
Next we define two functions, $h_{T}(x):\mathbb{R}^{d}\rightarrow\mathbb{R}$ and $h_{0}(x):\mathbb{R}^{d}\rightarrow\mathbb{R}$, as following
$h_{T}(x)$ =$=\begin{cases}\exp\left(-\frac{1}{2}x^{\prime}A_{T}x\right)2\pi\left(\theta^{*}\right)&\text{ if }|x|<\varepsilon^{*}\sqrt{T},\\ e^{-T\delta/2}\pi\left(\theta_{T}+x/\sqrt{T}\right)&\text{ if }|x|\geq\epsilon^{*}\sqrt{T}\end{cases}$
$h_{0}(x)$ $=\exp\left(-\frac{1}{2}x^{\prime}A_{0}x\right)2\pi\left(\theta^{*}\right)$
where $A_{T}^{a}=\operatorname{\Gamma}_{T}^{a}-\alpha^{a}I_{d}$, $A_{0}^{a}=\operatorname{\Gamma}_{*}^{a}-\alpha^{a}I_{d}$, $A_{T}=\sum_{a}\frac{|T_{a}|}{T}A_{T}^{a}$ and $A_{0}=\sum_{a}w_{a}A_{0}^{a}$. We note that, if
  1. (a)
    $g_{T}\rightarrow g_{0}$ and $h_{T}\rightarrow h_{0}$ pointwise,
  2. (b)
    $\int h_{T}\rightarrow\int h_{0}$,
  3. (c)
    $g_{T}=\left|g_{T}\right|\leq h_{T}$ for all $T$ sufficiently large, and
  4. (d)
    $g_{T},g_{0},h_{T},h_{0}\in L^{1}$ for all $T$ sufficiently large.
then by the generalized dominated convergence theorem, $\int g_{T}\rightarrow\int g_{0}$ and $\int\left|g_{T}-g_{0}\right|\rightarrow 0$ (e.g., Folland, 2013, exercises 2.20, 2.21). Since $\int q_{T}=1$, by equation (A.1) we have

### v2

@@ 3.3 Sequential Monte Carlo with gradient-based kernel
     Evaluate $\widehat{\nabla_{\theta}}{U}(\bm{\theta}_{(t-1)})$
     $\mathbf{\kappa}_{(t)}=\mathbf{\kappa}_{(t-1)}-(\bm{\eta}_{(t-1)}\mathbf{\kappa}_{(t-1)}+\widehat{\nabla_{\theta}}{U}(\bm{\theta}_{(t-1)})G(\bm{\theta}_{(t-1)}))\epsilon+\mathcal{N}(0,2a\epsilon G(\bm{\theta}_{(t-1)}))$
     $\bm{\theta}_{(t)}=\bm{\theta}_{(t-1)}+\mathbf{\kappa}_{(t)}G(\bm{\theta}_{(t-1)})\epsilon$
     $\bm{\eta}_{(t)}=\bm{\eta}_{(t-1)}+(\frac{1}{p}\mathbf{\kappa}_{(t)}^{T}\mathbf{\kappa}_{(t)}-1)\epsilon$
@@
When the target distribution is the scoring rule posterior, $U(\theta)=S(P_{\theta},\mathbf{y})$, where $P_{\theta}$ is a simulator model proposed for the observations $\mathbf{y}$. SRs such as the energy score can be expressed as $S(P_{\theta},\mathbf{y})=\mathbb{E}_{X,X^{\prime}\sim P_{\theta}}g(X,X^{\prime},\mathbf{y})$. Moreover, a simulation from the model $P_{\theta}$ can also be represented as $x=h_{\theta}(z)$, with $z\sim Z$, where the distribution $Z$ is independent of $\theta$. Pacchiardi et al. (2024b) showed that, if both $g$ and $h_{\theta}$ are differentiable, then an interchange of expectation and gradient step produces an unbiased estimate of the gradient $\widehat{\nabla_{\theta}}U(\theta_{u})$ using some random draws of $z_{i}\sim Z$ for $i=1,2,\ldots m$. This property is particularly useful in high-dimensional parameter spaces, where efficient exploration requires gradient information.
@@ 4 ETS with generalized posterior
A detailed proof of the above theorem can be found in Appendix B. The theorem suggests that, regardless of the number of episodes observed, the sequence of policies obtained from a policy iteration method integrated with ETS progressively converges toward optimal behavior as the iterations increase. The second term $\zeta_{l}(k,n)$ involves the samples drawn from the posterior of the model parameters obtained after observing up to $k$ episodes of interaction with the environment. As the episode count $k$ and the number of posterior samples $n$ increase, this term shrinks to 00 for any $l=1,2,\ldots$ according to Theorem 5 as the posterior distribution of the model parameters concentrates around the expected prequential score minimizer.
@@ 5.1 Finite action MDP
To demonstrate the ETS algorithm, we use the ‘inverted pendulum’ task from Dimitrakakis and Tziortziotis (2013). Here the agent targets to keep the pendulum upright as long as possible by switching actions. The state of the environment is defined as (the angle, and angular velocity) of the pendulum. There are three possible actions from each state. The actions are the force (in Newtons) applied in a certain direction, the action space is $(+50,0,-50)$. The physical equation for the system has 6 parameters: the pendulum mass, the cart mass, the pendulum length, the gravity, the amount of uniform noise, and the simulation time interval. The true value of the parameters are $\theta_{0}=(2.0,8.0,0.5,9.8,10,0.01)$ which is unknown, however, the simulator model ($P_{\theta}$) for the dynamics of the environment is known to us. In this environment, the agent gets a $+1$ reward for every balancing step.
In our comparison, we evaluated the online performance of both the model-free and Bayesian model-based approaches in the pendulum domain. In the model-free approach, we updated our policy after each episode of interaction with the environment. The policy updates ceased once we reached the maximum reward, which is known to be $1000$ in our setup because an episode is defined by a maximum of 1000 steps or until the pendulum falls for the first time.
@@
For policy learning, we adopted the least-squares policy iteration (LSPI) method proposed by Lagoudakis and Parr (2003). LSPI is a model-free approximate policy iteration method that leverages the least squares temporal-difference learning algorithm to approximate the $Q$ function. Specifically, we approximated the $Q$ function using a linear combination of a $4\times 4$ grid of Gaussian radial basis functions in LSPI, with a learning rate of $\gamma=0.99$.
Additionally, for the Bayesian approach, we assumed a uniform prior over the interval $[0.5\theta_{0},5.0\theta_{0}]$ for the model parameters. We obtained samples from the posterior distribution after each episode using SMC with the gradient-based Markov kernel (Jones and Leimkuhler, 2011) (refer to Table 1 in Appendix C for details of the hyperparameters). To define the intermediate target distributions between consecutive episodes, we chose the sequence of temperatures $\{\alpha_{k,l}\}_{l=1}^{L}$ at the $k$-th episode, such that the effective sample size (ESS) declines uniformly throughout the SMC iterations. We used the bisection method to find the temperature, $\alpha_{k,l}$ such that $\text{ESS}_{l}=c_{0}\times\text{ESS}_{l-1}$, the parameter $c_{0}$ for each experiment is mentioned in Appendix C.1. Once the ESS drops below half the original sample size, the particles are resampled. Here we used the zeroth order gradient (discussed in Appendix C.2) to compute an estimate of the gradient of the prequential loss. We used LSPI integrated with ETS to learn an optimal policy, as described in Section 2.1 where we updated the policy for a maximum of $J=30$ iterations in between each episode.
Figure 4 illustrates that both the model-free and Bayesian model-based methods exhibit improved policy learning with increased training data. However, the Bayesian approach achieves optimal rewards much sooner compared to model-free training. This expedited convergence in Bayesian methods can be attributed to the fact that once the posterior distribution of the model parameters converges, simulated interactions closely resemble true interactions. With this concentrated posterior distribution, longer trajectories of simulated data can be generated, facilitating a more accurate estimation of the value function. Consequently, when using these longer chains of simulated data to estimate the value function and perform LSPI to find an optimal policy, convergence occurs more rapidly due to the utilization of a more stable estimate of the $Q$-function.
@@
To demonstrate the application of ETS for the case where the simulator model is not known, we revisit the ‘inverted pendulum’ experiment with the previously mentioned parameters. Conditional GANs (Charlesworth and Montana, 2020; Zhao et al., 2021) have been widely adopted in the literature to model the dynamics. We have used a generative neural network, skipping the adversarial training, to model the difference between the next state and the current state conditioned on the current state and action as suggested by Nagabandi et al. (2018) and Deisenroth et al. (2013). Further, for the states $s$ that represents the angle, we have considered feeding in $(\sin(s),\cos(s))$ as inputs to the model. The model can be written as,
@@
For parameter inference we have used Sequential Monte Carlo (SMC) (implementation details can be found in Appendix C.1). In high-dimensional parameter spaces, selecting a well-chosen initial sample from an informative prior is crucial for effective Bayesian inference. To do so, we ran the Adam optimizer (Kingma, 2014) with a learning rate of $0.001$ for $1000$ steps, saving the final $100$ points. The covariance of these optimized points was computed and used to add Gaussian noise, generating a total of $300$ particles to initialize the SMC. To define the intermediate target distributions between consecutive episodes, we chose the sequence of temperatures $\{\alpha_{k,l}\}_{l=1}^{L}$ at the $k$-th episode, such that the conditional effective sample size (CESS) (Zhou et al., 2016) stays constant throughout the SMC iterations. We used the bisection method to find the temperature, $\alpha_{k,l}$ such that $\text{ESS}_{l}=c_{0}\times N$, where $N$ is the sample size and the parameter $c_{0}$ for each experiment is mentioned in Appendix C.1. Once the ESS drops below $N/2$, the particles are resampled.
@@
Figure 5 demonstrates that LSPI combined with ETS learns the optimal policy significantly faster than its model-free counterpart. Although the true dynamics of the environment were unknown, the generative neural network effectively learned the dynamics, accelerating the policy learning process. Note that the Theorem 5 on the consistency of the generalized posterior distribution and the Theorem 6 on the bound on the approximation error of the $Q$ function under ETS assume an unimodal posterior for model parameters. While this assumption may not hold when using a generative neural network as a simulator model, we still demonstrate that ETS is more sample-efficient for simpler models.
@@ 5.2 Continuous action MDP
When the action space is continuous, calculating the $Q$-function for all possible state-action pairs is impossible. A deterministic policy that maximizes the $Q$-function over the entire action space cannot be easily found. In such cases, a parameterized probabilistic policy is used, where $\mu_{\alpha}(a|s)$ denotes the probability of taking action $a$ when the environment is in state $s$, and $\alpha\in\mathbb{R}^{d^{\prime}}$ are the policy parameters. Usually, the policy distribution is considered as a Gaussian distribution and $\alpha$ would contain the mean and covariance of the distribution. The agent here tries to update the policy parameters according to the rewards collected during real-time interaction with the environment.
@@
To demonstrate the ETS strategy on a problem with continuous action space for the case of a misspecified model, we choose the ‘Hopper’ experiment from the OpenAI Gymnasium package (Kwiatkowski et al., 2024). This problem involves moving a two-dimensional, single-legged structure forward by applying force to three joints connecting its four body parts. Actions here consist of torques (ranging from -1 Nm to 1 Nm) applied to each of the three joints, while the environment’s state is defined as a $12$-dimensional real-valued vector of joint angles, angular velocities, positions, and velocities of the body parts. The reward function incentivizes the forward movement of the hopper while penalizing the application of excessive torque, which could destabilize the system.
We have used a similar neural network architecture as described in Section 5.1, for our simulator model, using three fully connected layers with $20$ nodes each. For sampling from this posterior of the model parameters, we used SMC, as before, but to maintain computational efficiency we used only $45$ particles (this figure was chosen based on the runs with 100 and 500 particles, which did not provide any significant improvements while already being expensive) for this problem. As before, we ran the Adam optimizer (Kingma, 2014) with a learning rate of $0.001$ for $1000$ steps, saving the final $15$ points. The covariance of these optimized points was computed and used to add Gaussian noise, generating a total of $45$ particles to initialize the SMC. Similar to the misspecified ‘inverted pendulum’ problem, we defined intermediate target distributions for SMC based on the CESS.
For policy learning, we applied the ‘REINFORCE’ (Williams, 1992) a foundational policy gradient method suitable for continuous action spaces, integrating ETS via equation (15). The policy parameters were then updated based on equation (16). Full implementation details of the experiment can be found in Sections C.1 and C.2 in the Appendix. Given that with an increasing amount of data, the posterior distribution can become overly concentrated, we limited posterior updates and sampling to the first $15$ episodes. At episode $15$, we halted model training and proceeded with the standard REINFORCE updates to refine the policy during future interactions with the environment.
Figure 6: For visual clarity we have considered the moving average (lag 10) of episodic returns collected using model-free REINFORCE and ETS-integrated REINFORCE. The average of the returns over five different seeds is plotted for both cases.
@@ 6 Conclusion
In this work, we have introduced a robust Bayesian framework for model-based reinforcement learning. A fully Bayesian treatment of model parameters traditionally requires a tractable likelihood function of parameters. This is often unavailable when the environment’s dynamics are modeled through generative neural networks, but is easier to simulate from. To address this issue, we formulated a generalized posterior for the model parameters using prequential scoring rules based on a Markovian assumption on observed trajectories, enabling generalized Bayesian inference in the absence of a known likelihood. Additionally, we established a Bernstein-von Mises (BvM) type consistency result for the proposed prequential scoring rule posterior in the discrete action setting, leaving the proof for the continuous action space for future work.
@@
For efficient sampling from the generalized posterior, we use SMC samplers. We use an adjusted SGLD kernel as the forward kernel in SMC, which handles noisy gradient estimates of the potential. To further improve sampling efficiency, we use gradient-based preconditioning, similar to the Adam optimizer, to better guide particles toward high-probability regions of the parameter space. For policy learning, we extended the classical TS by incorporating full posterior samples for enhanced policy search, introducing the expected Thompson sampling (ETS) approach. Using the BvM result, we derived an error bound to approximate the $Q$-function when a policy iteration method is integrated with ETS in well-specified model settings.
@@
To empirically evaluate the proposed method, we first compared classical model-free LSPI with the ETS-integrated Bayesian version on a simple inverted pendulum-balancing task. In both well-specified and misspecified model settings, the ETS-integrated approach learned the optimal policy more quickly than the model-free baseline. We then applied ETS to the more complex problem of teaching a single-legged hopper to move forward without falling. Even with continuous action space and a misspecified model, the ETS-based approach discovered a high-reward policy significantly faster than the model-free approach.
@@
In conclusion, we present a robust framework for policy learning that enables rapid identification of high-reward policies in complex tasks. This approach is especially relevant for applications in the design of clinical trials, robotics, and autonomous systems where agents must make quick decisions without the luxury of prolonged learning periods. However, we note that, due to the use of deep generative neural networks as a model for the MDP underlying the environment, the posterior of model parameters can be multimodal. This makes sampling difficult, especially in high dimensions. This may be dealt with using more informative priors (e.g., shrinkage priors) on the parameter space, which we keep as a future direction to explore. Another promising avenue could be through imposing a posterior distribution on the parameters of the policy in addition to the model parameters, hence learning the model parameters and optimizing the policy concurrently.
Appendix A Proofs related to the asymptotic behaviour of generalized posterior
A.1 Proof of Lemma 1
Lemma 1Under assumption A1, there exists a function $\operatorname{\mathcal{PS}}^{*}(\theta)$ such that as $T\to\infty$, $\frac{1}{T}\widetilde{\operatorname{\mathcal{PS}}}_{T}(\theta)\to\operatorname{\mathcal{PS}}^{*}(\theta)$ uniformly with probability one under $\mathbb{M}_{0}$.
Proof  We can define a statistical divergence between the proposed model $\mathbb{M}_{\theta}$ and the true distribution $\mathbb{M}_{0}$ in terms of the expected prequential scores upto time $T$ when the associated scoring rule is strictly proper (by assumption A1). We denote it by
$D_{T}(\theta)=\widetilde{\operatorname{\mathcal{PS}}}_{T}(\mathbb{M}_{\theta},\mathbb{M}_{0})-\widetilde{\operatorname{\mathcal{PS}}}_{T}(\mathbb{M}_{0},\mathbb{M}_{0}),$
where $\widetilde{\operatorname{\mathcal{PS}}}_{T}(\mathbb{M}_{0},\mathbb{M}_{0})$ is the generalized entropy associated with the model $\mathbb{M}_{0}$. Since the divergence $D_{T}(\theta)$ is of order $T$ (Dawid and Musio, 2014), $\frac{1}{T}D_{T}(\theta)$ has a finite limit and we assume that,
$\lim_{T\to\infty}\frac{1}{T}D_{T}(\theta)=D^{*}(\theta).$
Again from assumption A1 we have $\frac{1}{T}\widetilde{\operatorname{\mathcal{PS}}}_{T}(\mathbb{M}_{0},\mathbb{M}_{0})$ converging to a constant $c^{*}$ (say) as $T\to\infty$. Then,